OpenStudy (anonymous):

How would I solve this? -9=6/y Solve for y. ????

OpenStudy (mertsj):

Think of it like this:

OpenStudy (mertsj):

\[\frac{-9}{1}=\frac{6}{y}\]

OpenStudy (mertsj):

and cross multiply

OpenStudy (anonymous):

multiply both sides by y then divide by -9 on both sides

OpenStudy (anonymous):

So -9y/6?

OpenStudy (anonymous):

no, when you multiply by y, you end up with -9y = 6

OpenStudy (mertsj):

-9y=6

OpenStudy (anonymous):

when you divide by -9 you get y = -6/9 or -1/3

OpenStudy (mertsj):

-6/9 IS NOT -1/3

OpenStudy (anonymous):

my bad, thinking too fast haha.

OpenStudy (mertsj):

Happens to all of us.

OpenStudy (anonymous):

So once we get to -9y=6 what do we do from there?

OpenStudy (mertsj):

Divide both sides by -9

OpenStudy (anonymous):

-9/-9=1

OpenStudy (mertsj):



OpenStudy (mertsj):

The two sides are -9y and 6. =9y is the left side. 6 is the right side.

OpenStudy (mertsj):

Divide both sides by -9

OpenStudy (mertsj):